OpenStudy (mateeldaa):

Solve the equation in the interval [0, 360). csc theta = 1+ cot theta

OpenStudy (anonymous):

let me replace theta with x for ease of typing csc(x) = 1 + cot(x) 1/sin(x) = 1 + cos(x)/sin(x) by definition 1 = sin(x) + cos(x) by multiplying both sides by sin(x) 1 = sqrt(2) * sin(x + 45) by auxillary angle method 1/sqrt(2) = sin(x + 45) we know that sin(45) = 1/sqrt(2), and also sin(90 + 45) = 1/sqrt(2) also sin(360 + 45) = 1/sqrt(2) lets see... sin(x + 45) = sin(45), sin(90 + 45), sin(360 + 45) therefore x = 0, 90, 360 however, you said the interval does not include 360...therefore reject that value also notice that we had 1/sin(x) which means that sin(x) cannot equal 0. however at x = 0, sin(x) = 0, but not at x = 90. Therefore the only solution is x = 90 degrees

OpenStudy (anonymous):

Or in your case theta

OpenStudy (mateeldaa):

Thank you so much for your help! :)

OpenStudy (anonymous):

Anytime